### v1

# 0.1 Simple Estimation of Bi-Lipschitz Constants at Initialization

See Appendix LABEL:ap:exp_setup for further details on the setup of each experiment.
0.1 Simple Estimation of Bi-Lipschitz Constants at Initialization
First of all, we verify that BLNN behaves as expected with different values of $\alpha$ and $\beta$. We set all $g_{i}$ as ReLU or softplus functions. The Legendre-Fenchel transformation step is executed with GD (since it is simple and has good theoretical guarantees) until 1000 iterations are reached or the stopping condition, i.e., $\|\nabla_{y}\left(\langle y,x\rangle-F(y)\right)\|<10^{-3}$, is satisfied. The BLNN has input dimension 2, and $\alpha$ is set to 4 as varying it does not severely impact the fundamental quality of the theoretical bounds. On the other hand, $\beta$ was changed following the sequence $\{0.05+99.95\cdot j/100\}_{j=0}^{100}$. See Figures 1 and 2 for results.
Figure 1: Estimated Lipschitz and inverse Lipschitz constants of BLNN with gradient descent with softplus activation functions. The left figure is with 3 hidden layers and the right with 10. The $x$ axis corresponds to $j$ with $\beta=0.05+99.95\cdot j/100$.
Figure 2: Estimated Lipschitz and inverse Lipschitz constants of BLNN with gradient descent with ReLU activation functions. The left figure is with 3 hidden layers and the right with 10. The $x$ axis corresponds to $j$ with $\beta=0.05+99.95\cdot j/100$
As we can observe, our model combined with gradient descent can generate functions with bi-Lipschitz constants that are inside the predefined bounds as expected for both softplus and ReLU activation functions. That is why, it seems reasonable to use the simple gradient descent in order to execute the Legendre-Fenchel transformation.
Theoretically, since the initialization of weights are executed at random, the Lipschitz and inverse Lipschitz constants can be any value between $\alpha$ and $\alpha+\beta$. Curiously though, the estimated Lipschitzness is close to the upper bound for almost every model. Note that we are generating the network and then directly estimating the bi-Lipschitz constants, which means this phenomenon is possibly closely related to the initialization of weights. To better understand this phenomenon, consider the linear network $x\mapsto Wx$ where $W\in\mathbb{R}^{n\times n}$. In our architecture, this can be regarded as a simplification of $\nabla F_{\theta}$, where $F_{\theta}$ is the core ICNN. If we add $x/\beta$ to this network and compute the inverse operation, corresponding to $\nabla F_{\theta}^{\ast}$, we obtain $y\mapsto(W+I/\beta)^{-1}y$. By letting $\sigma_{\mathrm{min}}$ the smallest singular value of $W$, the Lipschitz constant becomes $(\sigma_{\mathrm{min}}+1/\beta)^{-1}$. As a result, if the initialization provides a $W$ with small $\sigma_{\mathrm{min}}$ close to 0, we inevitably obtain a function with Lipschitz constant close to the upper bound $\beta$, independently of the distribution of the other eigenvalues. This seems to happen in Figures 1 and 2. In other words, the construction employing the default initialization is biased, or more concretely, the derivative of an ICNN $F_{\theta}$ with initialized weights is with high probability $\epsilon$-inverse Lipschitz with $\epsilon<<1$. This is supported by Figure 3 which calculated the bi-Lipschitz constants of a $(4,60)$-BLNN initialized over 100 different trials. While the inverse Lipschitz constant is moderately distributed, the Lipschitz constant is concentrated on the theoretical maximum, i.e., 64.
We also modified the initialization procedure of the non-negative weights $W_{i}^{(z)}$ to confirm our hypothesis. The default approach was to draw each element from a uniform distribution proposed by X2010understanding and clamp all negative elements to 0. Figure 4 replaces this to a uniform distribution on an interval of $[0,1]$, and Figure 5 to $[1.0,1.1]$.
Figure 3: Histograms of Lipschitz and inverse Lipschitz constants of 100 $(4,60)$-BLNNs with non-negative weights $W_{i}^{(z)}$ initialized following the uniform distribution of X2010understanding. Negative elements were clamped to 0.
Figure 4: Histograms of Lipschitz and inverse Lipschitz constants of 100 $(4,60)$-BLNNs with non-negative weights $W_{i}^{(z)}$ initialized following a uniform distribution over $[0,1.0]$.
Figure 5: Histograms of Lipschitz and inverse Lipschitz constants of 100 $(4,60)$-BLNNs with non-negative weights $W_{i}^{(z)}$ initialized following a uniform distribution over $[1.0,1.1]$.
As we can conclude, the choice of the initialization impacts the distribution of the starting bi-Lipschitz constants for a $(\alpha,\beta)$-BLNN while theoretically any values are possible. In the default setting, we have a strong bias for the Lipschitz constant. Nevertheless, the fact that the bi-Lipschitz constants are close to the bounds can be regarded as the representation of a function with rich features. For example, if a matrix has the greatest eigenvalue and the smallest eigenvalue close to the limits, it implies that the other eigenvalues can assume any values in-between, allowing for considerable flexibility. This is a direct consequence of our meticulous construction of bi-Lipschitz neural networks respecting the desired constraints using convex neural networks and Legendre-Fenchel transformation.
The adequate initialization of the weights of an ICNN in order to obtain BLNNs with a richer distribution of bi-Lipschitz constants is a whole new problem and outside the scope of our work. We will not investigate further but it seems to be an interesting research topic since an appropriate initialization may facilitate training. In this paper, we used the original setup.
0.2 Tightness of the Bounds: Underestimation Case
This subsection provides additional results of experiment LABEL:exp:tight of the main paper. As a reminder, this experiment verifies the tightness of the proposed architecture by fitting the following function that has a discontinuity at the point $x=0$:
$f(x)=\left\{\begin{array}[]{ll}x&(x\leq 0)\\ x+1&(x>0)\end{array}\right.$
Therefore, when constraining the Lipschitzness of the model by a positive finite $L$, a model with tight Lipschitz bounds should achieve that upper bound around $x=0$ in order to reproduce the behavior of $f$. We compare our model with other Lipschitz constrained architectures that were introduced in Appendix LABEL:ap:rw. We take as comparison the spectral normalization (SN) (M2018spectral), AOL (P2022almost), Orthogonal (T2021orthogonalizing), SLL (A2023unified) and Sandwich (W2023direct). They can all regulate the Lipschitzness by an upper bound, but it is executed on a layer-wise basis.
Models are built so that $L$ is set to 2, 5, 10 and 50. For the compared prior works, this is realized by making all layers 1-Lipschitz, and scaling the input by $L$, which is the strategy often employed in practice. After learning was completed, the empirical Lipschitz constant was computed. The percentage between the empirical Lipschitz constant and the true upper bound $L$ can be found in Table 1. In order to be able to legitimately compare results, the number of parameters for each model were all of the same order.
[…]
(Nbr. of param.) $L=2$ $L=5$ $L=10$ $L=50$
SN (4.3K) 96.8%$\pm$1.3 80.1%$\pm$5.0 68.9 %$\pm$5.4 32.5%$\pm$5.8
AOL (4.3K) 96.2%$\pm$0.7 65.9%$\pm$4.7 46.5%$\pm$4.4 15.4%$\pm$2.8
Orthogonal (4.3K) 98.7%$\pm$0.4 75.1%$\pm$12.0 48.9%$\pm$11.7 14.6%$\pm$3.7
SLL (4.2K) 97.5 %$\pm$0.4 77.6%$\pm$6.5 50.0%$\pm$8.5 15.7%$\pm$6.5
Sandwich (4.5K) 97.5%$\pm$0.7 84.3%$\pm$2.0 60.2%$\pm$4.1 16.4%$\pm$3.6
Ours (4.3K) 100.2%$\pm$0.3 100.0%$\pm$0.0 99.4%$\pm$1.2 99.9%$\pm$0.1
Table 1: Tightness of Lipschitz bound with different methods. Each model was built with an upper-bound constraint on the Lipschitzness $L$, and the true Lipschitzness of the model after training was evaluated. The percentage between this value and $L$ is reported in the table, with a mean and standard deviation over five different seeds.
The difference in tightness is even clearer when setting the upper bound to 50 and plotting the result as shown in Figure 9. Only our method can capture the presence of a discontinuous jump around 0. We can conclude that our method clearly achieves the tightest bound for Lipschitzness as well as the most coherent fit of the function.
{subfigure}
[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption
Figure 6: SLL
Figure 7: Sandwich
Figure 8: Ours
Figure 9: Results of fitting the curve with SLL (first row), Sandwich (second row) and our method (third row) with a specified Lispchitzness of 50. The right column is a zoom of the right figure.
0.3 Flexibility of the Model: Overestimation Case
This subsection provides additional results of experiment LABEL:exp:flex of the main paper. As a reminder, this experiment verifies the flexibility of our model. While we already provided a theoretical guarantee of the expressive power of our model, it is important to show that it can learn without problem functions with smaller Lipschitz and larger inverse Lipschitz constants than those we set as parameters. At the same time, this addresses one concern of the previous experiment since the tightness of our model may be interpreted as a result of the strong influence the regularization terms we add, effectively generating only $\alpha+\beta$-Lipschitz functions. Therefore, we lead this experiment, where we overestimated the Lipschitz constant of the target function. The Lipschitz constant was set to $50$ but the function we want to learn is a linear function with slope 1. Experimental setting was the same as the previous subsection. Results are shown in Figures 13 and 17.
{subfigure}
0.9 Refer to caption Refer to caption {subfigure}0.9 Refer to caption Refer to caption {subfigure}0.9 Refer to caption Refer to caption
Figure 10: SN
Figure 11: AOL
Figure 12: Orthogonal
Figure 13: Results of fitting a linear function ($y=50x$ for left column and $y=x$ for right column) with SN (first row), AOL (second row) and Orthogonal (third row) with a specified Lipschitzness of 50.
{subfigure}
[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption
Figure 14: SLL
Figure 15: Sandwich
Figure 16: Ours
Figure 17: Results of fitting a linear function ($y=50x$ for left column and $y=x$ for right column) with SLL (first), Sandwich (second row) and our method (third row) with a specified Lipschitzness of 50.
Our model can without any problem fit the curve $y=50x$ and $y=x$ even in the overestimation scenario. Interestingly, this is not the case for prior work. On the one hand, SN and AOL cannot learn the linear function with slope 50. On the other hand, others cannot learn the simple identity function even after convergence. While the overall trend is along $y=x$, we can observe small fluctuations. This may be due to the large scaling factor $L$ we multiply at the input, leading to overfitting or interpolation which usually depends on the smoothness. That is, with high Lipschitzness, the function is likely to dynamically fluctuate around unseen points during training.
In order to examine the influence of this scaling factor on the training we run the experiment again with even larger $L$, namely, 100 and 1000. The evolution of the loss and the fitted function for $L=1000$ can be found in Figures 21 and 25, respectively. As we can observe, only our method is hardly affected by the size of $L$ and achieves a good generalization performance. While we do not have a rigorous mathematical explanation for this phenomenon, we believe our approach provides a better regularization than the previous methods. In other words, these layer-wise methods seem to provide a loss landscape that has many spurious minima where the function is driven. This phenomenon is aggravated with higher $L$. The difference in the smooth decrease of the loss function also supports this. The presence of these fluctuations due to the large scaling factor is thought to be related with the way we incorporated $L$ into the model. For previous work, we had no choice but to scale the function by multiplying with $L$. However, our method is adding a regularization term without touching the core function. This is an interesting avenue for future research.
In short, our model can perfectly learn both the function with slope 50 and 1 even though we greatly overestimated the Lipschitz constant.
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 18: AOL
Figure 19: Sandwich
Figure 20: Ours
Figure 21: The evolution of the loss with different upper bound constraints on the Lipschitz constant with AOL (top left), Sandwich (top right) and our method (bottom).
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 22: AOL
Figure 23: Sandwich
Figure 24: Ours
Figure 25: Results of fitting the linear function $y=x$ with AOL (top left), Sandwich (top right) and our method (bottom) with a specified Lipschitzness of 1000.
0.4 Summary of the Two Previous Experiments
This experiment summarizes the two previous experiments. We propose to run a regression task of the function $y=50x$ with SN (representing the layer-wise methods) and our model, where the Lipschitz bound $L$ is changed from $25$ to $125$. If the model provides tight bounds and perfect expressive power, then the loss should equal 0 for $L\geq 50$ as theoretically $y=50x$ can be reconstructed, and should increase once $L$ becomes smaller than 50 as the maximal slope we can reproduce is limited to $L$. Moreover, we expect that the optimization should proceed faster when $L$ is near $50$ as the initialization of the function should also be close to $y=50x$. Results are shown in Figure 28.
{subfigure}
[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption
Figure 26: SN
Figure 27: Ours
Figure 28: Results of fitting the linear function $y=50x$ with different $L$ for SN and our model. The left figure is the loss in function of $L$, and the right figure is the first iteration that the loss was below $0.5$ for each $L$ (if there is no point, it means that a loss below $0.5$ was never reached for this value of $L$). The red line emphasizes the point of $L=50$.
As we can observe, SN, a layer-wise model known to provide conservative bounds, only achieves a 0 loss from around $L=100$, while ours exactly exhibits an increase of the loss from $L=50$. This clearly shows that a layer-wise model has an expressive power that is largely lower than expected in the initial works. Notably, the convergence speed drastically decreases around $L=50$ for our approach. This suggests choosing bi-Lipschitz constants as tight as possible tailored to the specific problem to ensure faster convergence.
Interestingly, this difference of behavior around the true Lipschitz constant of our model could provide a new method to estimate the bi-Lipschitz constants. Indeed, by starting with a high Lipschitz constant $\alpha+\beta$ and decreasing it, we can find the largest point where the loss starts to increase. This value can become a good estimation of the Lipschitzness of the true function. We can similarly proceed for inverse Lipschitzness. We leave further investigation for future work.
0.5 Better Control for Annealing
Since our model provide tight bounds and good regularization performance as shown in the precedent experiments, the annealing of its parameters could be effective in some suitable settings. Let us consider the case where we want to learn the exponential function. It is not Lipschitz in general since its slope is monotonically diverging to $\infty$ as $x\to\infty$ and not inverse Lipschitz either since $e^{x}\to 0$ as $x\to-\infty$. We will start by a low Lipschitz parameter of the model and increase it by evaluating its effective Lipschitzness during training. The annealing will proceed in an elastic way. If the estimation is close to the upper bound then we will relax the bound since it means that our bound is too strong. That way, if we have a tight bound, the annealing process should progress faster and the overall function should rapidly converge to $\exp(x)$. The inverse Lipschitz constant is set to 0 and kept fixed. We compared our model with the Sandwich model. See Figure 31. We can verify that the annealing process is faster than the layer-wise model. Ours constantly wants to increase the Lipschitz constant while the Sandwich layer abandons at some point and cannot keep up the increase of the Lipschitzness.
{subfigure}
[b]0.9 Refer to caption Refer to caption {subfigure}[b]0.9 Refer to caption Refer to caption
Figure 29: Sandwich
Figure 30: Ours
Figure 31: Results of fitting $y=\exp(x)$ with Sandwich (top row) and our method (bottom row). The right figures are the evolution of the upper bound and the estimated Lipschitz constant over the iteration, and the left column the learned function. The $x$-axis of the right pictures represents the iteration number.
0.6 Uncertainty Estimation
0.6.1 Two Moons
This subsection provides additional results of experiment LABEL:exp:twomoons of the main paper. As a reminder, we compare the uncertainty when learning the two moons dataset with not only DUQ and DUQ+BLNN but also the Deep Ensembles method and a DUQ with no bi-Lipschitz regularization. Results are plotted in Figure 36. Uncertainty is calculated as the distance to the closest centroid (see Appendix LABEL:ap:unest for the mathematical formulation). Blue indicates high uncertainty, and yellow low uncertainty.
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 32: Deep ensemble
Figure 33: DUQ no reg.
Figure 34: DUQ
Figure 35: DUQ+BLNN
Figure 36: Uncertainty estimation with the two moons data set. Left figure is with a simple neural network without any constraints, and the right is with our model constraining bi-Lipschitzness.
$\alpha\backslash\beta$ 0.1 1.0 2.0 3.0 4.0 5.0
0.0 0.716$\pm$.111 0.874$\pm$.006 0.904$\pm$.045 0.990$\pm$.002 0.972$\pm$.043 0.957$\pm$.048
1.0 0.879$\pm$.004 0.991$\pm$.002 0.994$\pm$.004 0.992$\pm$.004 0.895$\pm$.195 0.993$\pm$.007
2.0 0.878$\pm$.006 0.990$\pm$.003 0.993$\pm$.004 0.936$\pm$.087 0.995$\pm$.003 0.518$\pm$.047
5.0 0.503$\pm$.006 0.500$\pm$.000 0.541$\pm$.080 0.503$\pm$.004 0.501$\pm$.002 0.503$\pm$.004
Table 2: Uncertainty quantification of two moons dataset with DUQ+BLNN with different $\alpha$ and $\beta$. Mean and standard deviation over five different seeds.
Figure 37: Uncertainty quantification of two moons dataset with DUQ+BLNN with a high $\alpha$. $\alpha=5.0$ and $\beta=3.0$. We do not show the points so that the highly certain area is visible.
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 38: DUQ+$(0,4)$-BLNN
Figure 39: DUQ+$(1,4)$-BLNN
Figure 40: DUQ+$(2,4)$-BLNN
Figure 41: Uncertainty quantification of two moons dataset with DUQ+BLNN with different $\alpha$.
As mentioned in the main paper, DUQ+BLNN achieves a tighter bound which is possible because we are able to tune the bi-Lipschitz constant unlike DUQ. The experimental setup for DUQ was chosen like the original paper (V2020uncertainty).
Table 2 provides details on the accuracy of the learned model with different $\alpha$ and $\beta$, used for the grid search. As for the accuracy, the value of $\alpha$ does not matter as long as $\beta$ is chosen large enough. Indeed, a BLNN with $(\alpha,\beta)=(0.0,3.0)$ theoretically includes that with $(\alpha,\beta)=(1.0,2.0)$. That is why they exhibit similar performance. Nevertheless, a too loose setting of these parameters, especially for $\beta$, still seems to affect the training, and a as tight as possible choice is beneficial as we can see for $(\alpha,\beta)=(2.0,5.0)$. In a sense, decreasing $\beta$ which is mainly in charge of the Lipschitz constant helps better generalization as pointed out in prior work as well (B2017spectrally). In the same direction, a too large value of $\alpha$ leads to a nonsensical result as we can understand when $\alpha=5$. In this case, the model is randomly guessing which means that it is unsure everywhere. This is also translated in the uncertainty estimation as shown in Figure 37. The area where the model is unsure increases as $\alpha$ is increased, meaning that it becomes more sensitive to unknown data (see Figure 41). Comparatively, setting $\alpha=0$ does not affect so much since the injectivity constraint is already beneficial in this type of task. Finally, a too low value of $\beta$ means that the inverse Lipschitz and Lipschitz constants are close to each other and the function behaves more like a linear function. Consequently, learning becomes difficult as we can conclude from the table with $\beta=0.1$.
0.6.2 FashionMNIST
This subsection provides additional results of experiment LABEL:exp:fashion of the main paper. In this experiment, we use real world data of FashionMNIST (X2017fashion), MNIST (L2010mnist) and NotMNIST (B2011notmnist). The task is to learn to classify FashionMNIST but at the same time to detect out-of-distribution points from MNIST and NotMNIST. This task to distinguish FashionMNIST from MNIST datasets is known to be a complicated task (V2020uncertainty). We compute the AUROC for the detection performance. Results for a downsampled dataset and full size dataset are shown in Table 3 and 4, respectively. A visualization of the ROC curve can be found in Figures 44 and 47.
[…]
NotMNIST
DUQ 0.8564$\pm$.0066 0.078$\pm$.004 0.870$\pm$.040 0.852$\pm$.010
DUQ+BLNN 0.8595$\pm$.0032 0.078$\pm$.004 0.876$\pm$.008 0.960$\pm$.009
Table 3: Out-of-distribution detection task with down-sampled data. FashionMNIST vs MNIST and FashionMNIST vs NotMNIST dataset with DUQ and DUQ+BLNN. For the BLNN, $\alpha=0.2$ and $\beta=0.4$. Mean and standard deviation over five trials.
[…]
NotMNIST
DUQ 0.8893$\pm$.0037 0.064$\pm$.005 0.862$\pm$.035 0.900$\pm$.024
DUQ+BLNN 0.8985$\pm$.0040 0.060$\pm$.000 0.896$\pm$.008 0.964$\pm$.005
Table 4: Out-of-distribution detection task with full size data. FashionMNIST vs MNIST and FashionMNIST vs NotMNIST dataset with DUQ and DUQ+BLNN. For the BLNN, $\alpha=0$ and $\beta=3.0$. Mean and standard deviation over five trials.
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 42: FashionMNIST vs MNIST
Figure 43: FashionMNIST vs NotMNIST
Figure 44: ROC between dwonsampled FashionMNIST vs MNIST (left) and FashionMNIST vs NotMNIST (right).
{subfigure}
[b]0.45 Refer to caption {subfigure}[b]0.45 Refer to caption
Figure 45: FashionMNIST vs MNIST
Figure 46: FashionMNIST vs NotMNIST
Figure 47: ROC between full size FashionMNIST vs MNIST (left) and FashionMNIST vs NotMNIST (right).